Grade 12th passMechanics

A particle is projected upward to a maximum height H. find the time after which the ratio of potential energy to kinetic energy is 3:1

Sanju Kumar
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear student
 
 
The total energy given to the ball when thrown is equal to mgh
then initially it was only \frac{1}{2}mv^{2}
 
Now by conservation of energy
\frac{1}{2}mv^{2}=mgh=Total Energy…........................(1)
 
as KE + PE = total energy
so when the ratio is 1:3
K.E. = \frac{1}{4} T.E.
TE  means total energy 
 
After time t,
velocity = v-gt
 
\frac{1}{2}m(v-gt)^{2} =\frac{1}{4}(\frac{1}{2} mv^{2})
 
Solving for t we get a equation in v.
Now substitute v in terms of H from equation (1) and get the answer...
 
Hope it helps